OpenStudy (anonymous):

Write the standard form or the equation. Center (6,8) R=10

OpenStudy (amoodarya):

\[(x-\alpha)^2 +(y-\beta)^2=r^2\\(\alpha,\beta)=(6,8) \\r=10\]

OpenStudy (amoodarya):